Try to find the answer to 3(y-7)=27 if u don't mind
Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

3y-21=27
3y=27+21
3y=48
y=48/3
y=16

Answer by Alan3354(69443)

Depends on the question.